OpenStudy (shreenithaa):

write the solution in interval notation using the union sign U 2y+5<5y-1 or 5y-1<3y-7

OpenStudy (shreenithaa):

@ganeshie8 @zepdrix

zepdrix (zepdrix):

\[\large\rm 2y+5\lt 5y-1\]What do you get when you simplify? :)

OpenStudy (shreenithaa):

2y+5<5y-1 or 5y-1<3y-7 -5 -5 +1 +1 2y<5y-6 or 5y<3y-6 -5y -5y -3y -3y -3y<6 or 2y<-8 //////////////// -3 2 y>-2 or y<-4

OpenStudy (shreenithaa):

@zepdrix

OpenStudy (shreenithaa):

how do u write the interval notation using the union sign

zepdrix (zepdrix):

Hmm, those don't look quite right. Here is what I would recomment for the first relationship.\[\large\rm 2y+5\lt 5y-1\]Subtracting 2y from each side,\[\large\rm 5\lt 3y-1\]Adding 1,\[\large\rm 6\lt 3y\]Divide 3 gives us,\[\large\rm 2\lt y\]

OpenStudy (shreenithaa):

how do u do the union sign?

zepdrix (zepdrix):

Again, to do it in text, you'll have to google it :) lol In the equation tool, it's the command \cup

OpenStudy (shreenithaa):

no i mean write it in interval notation

zepdrix (zepdrix):

Well we're not quite there yet :O Both of your y values look incorrect. The other y interval I think is \(\large\rm y\lt -3\) That's what I'm coming up with at least, hmm

OpenStudy (shreenithaa):

i checked and i got t right

zepdrix (zepdrix):

...?? 0_o -2 and -4 are not the correct values though..

OpenStudy (shreenithaa):

yeah -2 isnt \ just checked forgot it was -6

OpenStudy (shreenithaa):

-4 is right though

zepdrix (zepdrix):

No, neither number is correct :( One sec..

zepdrix (zepdrix):

Ya in the first one, it was supposed to be -6, not 6. So that should change it to \(\large\rm y\gt 2\) In the second one, your -6 turned into a -8 halfway through. So fixing that should give you \(\large\rm y\lt -3\)

OpenStudy (shreenithaa):

yeah nevermind

zepdrix (zepdrix):

Hard to keep track of all these little steps :D lol

OpenStudy (shreenithaa):

i c my mistake sorry

OpenStudy (shreenithaa):

now finally how do u do the union thingy

zepdrix (zepdrix):

\(\large\rm y\lt-3 ~OR~ y\gt2\) The word OR means union, as you said :)\[\large\rm y\lt-3 ~\cup~ y\gt2\]Let's write the y<3 as an interval.

zepdrix (zepdrix):

y<-3*

OpenStudy (shreenithaa):

so the u is replaced for or?

zepdrix (zepdrix):

Yes :) The union symbol and the word OR mean the same thing

zepdrix (zepdrix):

But we have a little more to do

OpenStudy (shreenithaa):

ooh wow does that sound easy now what grade r u in and state

OpenStudy (shreenithaa):

i only need the union sign

zepdrix (zepdrix):

\(\large\rm y\lt-3\) can be written as \(\Large\rm \left(-\infty,-3\right)\) -3 is the largest value this interval can have. and every number below that should be included in the interval. Does that make sense? :o

zepdrix (zepdrix):

oh you don't need the intervals? :D

OpenStudy (shreenithaa):

nope the intervals have to be as the union sign or something like that

OpenStudy (shreenithaa):

these are disjunctions so yeah

OpenStudy (shreenithaa):

lol

zepdrix (zepdrix):

If we're using the Union symbol, we usually write the inequalities as intervals. So for our problem,\[\large\rm y\lt-3 ~OR~ y\gt2\]We would write it as\[\large\rm (-\infty,-3)\cup(2,\infty)\]

OpenStudy (shreenithaa):

thnx what grade r u in